OpenStudy (calculusxy):

QUADRATIC Word-Problem Engineers are programming a robotic arm to move toward an object it will pick up. They define the way the arm will move by using the function f(x) = -1/2x^2 - 2x + 16, where f(x) represents the distance between the arm and the object, and x represents the number of seconds that arm is in motion. Determine the number of seconds it will take for the arm to reach the object it will pick-up.

OpenStudy (calculusxy):

@jhonyy9 @zepdrix

OpenStudy (calculusxy):

Does it mean for me to find the vertex?

zepdrix (zepdrix):

f(x) represents distance x represents time we want to know the `time x` when the `distance between the objects is zero, f(x)=0`

zepdrix (zepdrix):

\[\large\rm f(x) = -\frac12x^2 - 2x + 16\]\[\large\rm 0 = -\frac12x^2 - 2x + 16\]Solve for x, ya? :)

OpenStudy (calculusxy):

Okay!

zepdrix (zepdrix):

Ooo this one factors nicely :o

OpenStudy (calculusxy):

I used the quadratic equation.

zepdrix (zepdrix):

Fair enough :)

OpenStudy (calculusxy):

I got x = 8 , 4

zepdrix (zepdrix):

Hmm you're missing a minus sign on one of those values...

OpenStudy (calculusxy):

sorry -8, 4

zepdrix (zepdrix):

Niceeeee. So which value(s) make sense? At time -8 seconds, the arm has reached the object?

OpenStudy (calculusxy):

4

OpenStudy (calculusxy):

There's another part. It says: Describe how the engineers could modify the function so that it takes the arm one second longer to reach the object.

zepdrix (zepdrix):

hmm

zepdrix (zepdrix):

Well this is how I approached the problem,\[\large\rm 0 = -\frac12x^2 - 2x + 16\]Multiplying by -2,\[\large\rm 0 = x^2+4x-32\]Factors into,\[\large\rm 0=(x-4)(x+8)\]So if we want the arm to reach the object in 5 seconds instead of 4, we could replace this factor (x-4) with (x-5),\[\large\rm 0=(x-5)(x+8)\]And we would have to think about what that does to the original equation.

zepdrix (zepdrix):

That seems a little convoluted though, maybe that's not the approach they're looking for... hmm

OpenStudy (calculusxy):

If in the quadratic we have \(\frac{2 - 7}{-1}\), then it would be 5 seconds

OpenStudy (calculusxy):

We can change the c term to 18 I think.

zepdrix (zepdrix):

Ya you could do that maybe :) So you're assuming that the b value is still -2, and that the a is still -1/2. So we need the square root to give us 49.

zepdrix (zepdrix):

\[\large\rm 2^2-4\left(-\frac{1}{2}\right)c=49\]

zepdrix (zepdrix):

I think we get 22.5 or something.. hmm

jhonyy9 (jhonyy9):

@zepdrix i agree you - 5 need being a root of this quadratic so this mean that (x-5)(x-8) will result sure this quadratic

zepdrix (zepdrix):

There are many changes you could make to get a root of 5 though, so this route should be fine as well :) Did you follow what I was saying? In order to get that 7 we must have,\[\large\rm b^2-4ac=49\]

OpenStudy (calculusxy):

yes

zepdrix (zepdrix):

Ya I like your approach better xy, it's a much simpler way of altering the problem. You simply replace increase the 16 by 6.5 more to get the root x=5 that we're looking for.

zepdrix (zepdrix):

replace the 16 by 22.5* or increase the 16 by 6.5* blah my words got mixed up there.

OpenStudy (calculusxy):

Okay so the quadratic is \(f(x) = \frac{-1}{2}x^2 - 2x + 22.5\)

zepdrix (zepdrix):

Yes, good. Here it is graphed if you would like to verify the result. https://www.desmos.com/calculator/bahqvckilk

OpenStudy (calculusxy):

Okay thanks! Can you help me on one more question please?

zepdrix (zepdrix):

Sure!

OpenStudy (calculusxy):

An art teacher plans to spend a total of $100 on art supplies. The supplies can be purchased at a local art supply store or an online store. The actual teams are the same price in both the online store and local art shop. The Additional fees of sales tax, shipping, delivery, and processing vary between both stores.  Write two separate equations for each one of the stores with a total of $100. Use the equations to explain why the art teacher should or should not order through the local art supply shop.

OpenStudy (calculusxy):

My equation for the local art supply shop:\(1.07x + $5 = $100\) My equation for the online store: \(x + 0.2x = $100\)

zepdrix (zepdrix):

Ooo looks great so far!

OpenStudy (calculusxy):

So do you think that I should find the value of x for the second part to the whole question?

OpenStudy (calculusxy):

The value of x for both of the equations.